The volume V of a right circular cone is given by
V= 1/3[pi]r^2h.
The letter r stands for the radius of the circular base of the cone, and h is the height of the cone.
Solve for h.

Answer by stanbon(57282)

V= 1/3[pi]r^2h.
The letter r stands for the radius of the circular base of the cone, and h is the height of the cone.
Solve for h.
------------------
V= (1/3)[pi]r^2h
(pi)r^2h = 3V
h = [3V]/[(pi)r^2]
======================
Cheers,
Stan H.